Calculating CPI, real, and per capita terms.
Year: 2000, 2001, 2002.
Data: 847, 902, 911.
GDP ($B): 562, 607, 619.
Pop (m): 183, 201, 232.
Calculate GDP for 2002 using constant 2000 price terms
1
Expert's answer
2016-04-30T05:10:02-0400
Calculating CPI, real, and per capita terms.
Year: 2000, 2001, 2002.
Data: 847, 902, 911.
GDP ($B): 562, 607, 619.
Pop (m): 183, 201, 232.
CPI(2000) = 100,
CPI(2001) = 607/562*100 = 108,
CPI(2002) = 619/607*100 = 102.
GDP per capita (2000) = 562/183 = $3071,
GDP per capita (2000) = 607/201 = $3020,
GDP per capita (2000) = 619/232 = $2668.